OpenStudy (anonymous):

Leo was driving at 60km/ph. He wanted to increase his distance by 25% for the next part of the trip, be he also wanted to cut his time in half. How fast would he have to be going?

OpenStudy (anonymous):

He wants to go 60*1.25 = 75km in half an hour. s = 60*1.25*2 =150km/h

OpenStudy (anonymous):

That's the answer i got, but it took me 2 hours to get it, and i still wasn't sure. what is "s" ?? i was trying to d=r x t ????

OpenStudy (anonymous):

Wait... maybe i interpreted the q wrong.

OpenStudy (anonymous):

s = speed

OpenStudy (anonymous):

Re-re-reading the question i think my first answer is what they are asking.

OpenStudy (anonymous):

But he wants to go even FURTHER, 25%,,,,so he would have to go even FASTER then 120kph right??

OpenStudy (anonymous):

Yeah, it'll be 150km/h.

OpenStudy (anonymous):

Thanks SO much!!

OpenStudy (anonymous):

Your welcome. What made me second guess the question is that i was thinking .... "He wants to go 60*1.25 = 75km in half an hour." ... which isn't necessarly true. He just wants to increase his distance by 25%. So yeah the answer is definitely: speed = 60*1.25*2 =150km/h